# Exploring Multi-Document Information Consolidation for Scientific Sentiment Summarization

### v1

@@ Appendix D Prompt to Get Content and Sentiment Expressions with GPT-4
6Please first identify the sentence with judgements only on the quality of scientific manuscripts based on the criteria facets for scientific peer-review: novelty, soundness, clarity, advancement, compliance and overall quality within the given document.
@@ F.1 Prediction with Judgements of Source documents
The judgements are extracted from source documents, and they are in the same criteria facet to the target meta-review judgement.
@@ G.1 Prompt with Descriptive Consolidation Logic
92. Predicting Criteria Facets, Sentiment Levels, and Convincingness Levels;
10Candidate criteria facets: Novelty, Soundness, Clarity, Advancement, Compliance, and Overall quality
@@ 92. Predicting Criteria Facets, Sentiment Levels, and Convincingness Levels;
143. Reorganize extracted judgements in different clusters for different criteria facets;
@@
164. Generate a small summary for judgements on the same criteria facet with comparison and aggregation;
@@
185. Aggregate judgements in different criteria facets and write a meta-review based on the aggregation.
@@ G.2 Prompts Used in the Pipeline Generation
3Write a meta-review to summarize the above sub-summaries of reviews and discussions in different criteria facets for a manuscript.

### v2

@@ Appendix D Prompt to Get Content and Sentiment Expressions with GPT-4
6Please first identify the sentence with judgements only on the quality of scientific manuscripts based on the review facets for scientific peer-review: novelty, soundness, clarity, advancement, compliance and overall quality within the given document.
@@ F.1 Prediction with Judgements of Source documents
The judgements are extracted from source documents, and they are in the same review facet to the target meta-review judgement.
@@ G.1 Prompt with Descriptive Consolidation Logic
92. Predicting Review Facets, Sentiment Levels, and Convincingness Levels;
10Candidate review facets: Novelty, Soundness, Clarity, Advancement, Compliance, and Overall quality
@@ 92. Predicting Review Facets, Sentiment Levels, and Convincingness Levels;
143. Reorganize extracted judgements in different clusters for different review facets;
@@
164. Generate a small summary for judgements on the same review facet with comparison and aggregation;
@@
185. Aggregate judgements in different review facets and write a meta-review based on the aggregation.
@@ G.2 Prompts Used in the Pipeline Generation
3Write a meta-review to summarize the above sub-summaries of reviews and discussions in different review facets for a manuscript.
G.3 Prompts from Prompt-Naive
For Prompt-Naive in our experiments, the prompt we use is as follows.
1{{input_documents}}
2
3Write a meta-review based on the above reviews and discussions for a manuscript.
G.4 Prompts from Prompt-LLM
For Prompt-LLM, we have to generate first the steps with GPT-4 and then the meta-review based on the generated steps.
The prompt to generate the steps:
1{{input_documents}}
2
3What are the steps to write a meta-review specifically for the above reviews and discussions of a manuscript.
The prompt to generate the meta-review:
1{{input_documents}}
2
3Follow the following steps and write a meta-review based on the above reviews and discussions for a manuscript.
4
5{{generated_steps}}